I am a seven-digit number. My first five digits are 2, 1, 3, 5, and 8. I am divisible by 99. What are my last two digits?
I have tried 11, 23, 22, 25, 24, 69. I can't figure it out, will you help me?

Answer by richard1234(7193)

We can write the number in the form 21358ab, where a,b are digits.

We know that the sum of the digits is a multiple of 9, so 2+1+3+5+8+a+b ≡ 0 (mod 9) --> 19+a+b ≡ 0 (mod 9) --> a+b ≡ 8 (mod 9), a+b = 8 or 17.

An integer is a multiple of 11 if and only if the difference between the odd-positioned digits and the even-positioned digits is a multiple of 11. Hence, (2+3+8+b) - (1+5+a) ≡ 0 (mod 11), 7+(b-a) ≡ 0 (mod 11), b-a ≡ 4 (mod 11), a-b = 7 or -4.

We have the system of equations a+b = 8, 17 and a-b = 7, -4. If we add the equations, we expect an even number, so the only possible ordered pairs(a+b, a-b) are (8, -4) and (17, 7). Do each case separately:

a+b = 8
a-b = -4

This implies a = 2 and b = 6.

a+b = 17
a-b = 7

This implies a = 12, impossible.

The only possible case is a = 2, b = 6. It can be checked that 2135826/99 = 21574.